OpenStudy (anonymous):

What is the y-intercept of the quadratic function below? f(x) = 24x2 + 4x - 48 A. 6 B. -48 C. 0 D. 2

OpenStudy (marigirl):

its usually the constant (the number at the end), so in this case -48

OpenStudy (anonymous):

can you help with another??

OpenStudy (anonymous):

Please write your question here

OpenStudy (anonymous):

Identify the x-intercepts of the function below. f(x) = x2 + 9x + 20 A. 4 and 5 B. -4 and 5 C. -4 and -5 D. 0 and 20

OpenStudy (anonymous):

we have y = x^2 + 9x + 20 to find x -intercept put y =0 so we will get x^2 + 9x + 20 =0 then use quadratic formula to find values of x

OpenStudy (anonymous):

To find x intercept plug y = 0 or f ( x) = 0 into given equation can you try ?

OpenStudy (anonymous):

?? D?

OpenStudy (anonymous):

no please check it

OpenStudy (anonymous):

A!

OpenStudy (anonymous):

Please check again ..

OpenStudy (anonymous):

B

OpenStudy (anonymous):

nope .. :|

OpenStudy (anonymous):

C!

OpenStudy (anonymous):

Excellent .. :)

OpenStudy (anonymous):

I got A, B, and C confused!

OpenStudy (anonymous):

lol

OpenStudy (pinksapphire):

For the equation f(x) = 24x2 + 4x - 48 the right answer is C.0. try it on a scientific calculator and see where the line crosses the y-intercept.